During June, Danby Company’s material purchases amounted to 8,300 pounds at a price of $7.50 per pound. Actual costs incurred in the production of 1,700 units were as follows:

Direct labor: $ 142,695 ($18.90 per hour)
Direct material: $ 39,000 ($7.50 per pound)

The standards for one unit of Danby Company’s product are as follows:

Direct Labor: Direct Material:
Quantity, 4 hours per unit Quantity, 3 pounds per unit
Rate, $18.80 per hour Price, $7.20 per pound

Required:

Compute the direct-material price and quantity variances, the direct-material purchase price variance, and the direct-labor rate and efficiency variances. (Indicate the effect of each variance by selecting "Favorable" or "Unfavorable". Select "None" and enter "0" for no effect (i.e., zero variance).)

Direct-material price variance
Direct-material quantity variance   
Direct-material purchase price variance
Direct-labor rate variance
Direct-labor efficiency variance
Answer #1

SOLUTION

1. Direct-material price variance = (Actual price - Standard price) * Actual quantity

= ($7.50 - $7.20) * 5,200

= 1,560 U

Actual quantity = Actual direct material used / Actual material price per pound

= 39,000 / 7.50 = 5,200

2. Direct-material quantity variance = (Standard quantity - Actual quantity) * Standard price

= ((1,700 * 3) - 5,200) * 7.20

= (5,100 - 5,200) * 7.20

= 720 (U)

3. Direct-material purchase price variance = (Actual price - Standard price) * Purchased quantity

= ($7.50 - $7.20) * 8,300

= 2,490 (U)

4. Direct-labor rate variance = (Actual rate - Standard rate) * Actual hours

= (18.90 - 18.80) * 7,550

= 755 (U)

Actual hours = Actual direct labor cost / Actual labor cost per hour

= 142,695 / 18.90 = 7,550 hours

5. Direct-labor efficiency variance = (Standard hours - Actual hours) * Standard rate

= [(1,700 * 4) - 7,550] * 18.80

= (6,800 - 7,550) * 18.80

= 14,100 (U)